# Fast and Accurate Task Planning using Neuro-Symbolic Language Models and Multi-level Goal Decomposition

### v1

@@ Abstract
In robotic task planning, symbolic planners using rule-based representations like PDDL are effective but struggle with long-sequential tasks in complicated planning environments due to exponentially increasing search space. Recently, Large Language Models (LLMs) based on artificial neural networks have emerged as promising alternatives for autonomous robot task planning, offering faster inference and leveraging commonsense knowledge. However, they typically suffer from lower success rates. In this paper, to address the limitations of the current symbolic (slow speed) or LLM-based approaches (low accuracy), we propose a novel neuro-symbolic task planner that decomposes complex tasks into subgoals using LLM and carries out task planning for each subgoal using either symbolic or MCTS-based LLM planners, depending on the subgoal complexity. Generating subgoals helps reduce planning time and improve success rates by narrowing the overall search space and enabling LLMs to focus on smaller, more manageable tasks. Our method significantly reduces planning time while maintaining a competitive success rate, as demonstrated through experiments in different public task planning domains, as well as real-world and simulated robotics environments.
@@ I INTRODUCTION
At a high level, the use of LLMs for task planning is divided into treating LLMs as a policy model (known as L-Policy) or as a world model (known as L-Model)[4]. L-Policy exploits the commonsense knowledge of LLMs to directly query proper policy for a given state, while L-Model utilizes LLMs as a simulation model of the world to query the state of the world as a result of executing an action or a policy. However, despite their impressive ability, LLMs still struggle with issues like token inefficiency and correction inefficiency[5], generating hallucinatory action sequences and failing to identify accurate plans as task complexity increases[6]. To address the limitations of current LLM-based task planners, we propose a novel neuro-symbolic task planner that leverages LLMs as both L-Policy and L-Model to solve a long-sequential robotic task. Our planner is significantly faster than symbolic planners and more accurate than LLM-based planners.
An immediate issue in handling long-sequential tasks using LLMs is LLM’s token inefficiency since the planning descriptions involve a long and repetitive sequence of world and robotic states as well as a history of policies and their results. To circumvent this issue, we utilize LLMs as L-Model to generate a sequence of subgoals for a long-horizon task, effectively decomposing it into smaller and manageable sub-tasks. This goal decomposition also provides a useful side-effect to reduce the overall search space, yielding an accurate subgoal planner based on LLMs. Indeed, we use the Monte Carlo Tree Search (MCTS) algorithm while using LLMs as L-Policy to accurately solve each subgoal, reducing the correction inefficiency common in LLM-based planners. Furthermore, if the original task is moderately complex, requiring a smaller minimum description length (MDL)[4] to encode the given problem, one can rely on a symbolic planner to solve the subgoals exactly while effectively avoiding the exponential growth of planning time.
@@
    Planning formulation: Given a planning goal in natural language description and domain knowledge, our task planner relies on PDDL to encode the problem descriptions. We also obtain the semantic and spatial relationships of target objects in the environment using a multi-modal LLM, translated and encoded in problem PDDL.
@@
    Task planning: If the MDL is moderate, we rely on a symbolic planner to solve each subgoal; otherwise, we generate and expand a search tree and use the MCTS algorithm with L-Policy as a roll-out policy to solve the subgoal. This subgoal planning is repeated for each sub-task, and the plans are combined to form the overall plan.
We conducted experiments using a commercial LLM model across three task planning domains while varying the problem complexity. Compared to the state-of-the-art symbolic task planner, such as the fast-downward planner [3], our approach significantly reduced planning time while maintaining an acceptable success rate. Additionally, we conducted experiments using dual robot manipulators as well as using a robotic simulator to demonstrate the practical utility of our planner.
@@
The rest of this paper is organized as follows. In Sec. II, we review relevant work to task planning. In Sec. III, we outline the overall pipeline, and in Sec. IV, we explain the algorithms of both the symbolic subgoal planner and the LLM-based subgoal planner. In Sec. V, we present the task planning results and experiments in real and simulation robotics environments, and conclude the paper and discuss future work in Sec. VI.
@@ II-B LLM-based Robot Task Planning
Several recent studies have explored using LLMs for robot task planning, leveraging their understanding of the real world. [9] combines language understanding with action grounding in real-world affordances, enabling robots to execute tasks based on their capabilities. Similarly, [10] introduced a prompting scheme that enables LLM to generate Python codes composed of robot action primitives, incorporating environmental state feedback. TAMP has also been addressed using LLM by [11] and [12], enabling LLM as spatial relationship generators between environment objects. However, a common limitation of these approaches is low success rates in solving long sequential tasks.
@@ II-C Hybird Task Planning
Recently, studies have been conducted on integrating LLMs with symbolic planning methods. [13] and [14] used LLMs as translators between natural languages and symbolic languages like PDDL, converting natural language problem descriptions into PDDL initial states and goals through few-shot prompting. However, these approaches struggle in real-world applications where problems are not presented in natural language. [15] combined LLMs with vision models to generate planning problem specifications based on real-world scenes, using re-prompting to correct specification errors. Several studies have also explored using LLMs to solve PDDL problems. [16] showed that while LLMs can solve some non-trivial PDDL problems, they often fail on more complex tasks, though their outputs can guide heuristic planners. Building on this, [17] proposed a method where LLMs generate Python functions to create PDDL plans with automated debugging. [18] introduced a framework that iteratively refines PDDL plans using validator feedback. While these methods have improved success rates compared to LLM-only methods, they have been tested mostly on small-size problems.
@@ II-D Integrating LLMs with Tree Search
Combining tree structures with LLM-generated actions has been explored. [19] samples possible next actions from the current state using an LLM and selects the best action via an LLM-evaluator, repeating this process with DFS or BFS. To address this method’s token and runtime inefficiencies, [5] proposes sampling multiple plans at once rather than repeatedly calling the LLM to generate an action tree, and proposes selecting actions from the tree based on observations and histories. [20] integrates MCTS with LLMs by iteratively calculating state transitions in MDPs, and [4] also uses LLMs as both L-Model and L-Policy combined with MCTS to solve large-scale POMDPs.
Our method also uses an LLM to sample multiple plans, and exploit MCTS to find a plan. Still, it differs from [5] in that ours relies on LLM-induced goal decomposition to generate multiple subplans and solve a deterministic problem, unlike [4]. Our MCTS is performed on a fixed tree for a sub-problem, not the entire planning problem.
@@ III TASK PLANNING PIPELINE
$P\equiv\langle\mathcal{S},\mathcal{A},\mathcal{T},s_{0},S^{\star}\rangle,$ (1)
where $\mathcal{S}$ is a finite set of fully observable states, $\mathcal{A}$ is a finite set of possible actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is a deterministic state transition function, $s_{0}\in\mathcal{S}$ is an initial state, and $S^{\star}\subset\mathcal{S}$ is a set of goal states. Our planning objective is to find a policy $\pi=\{a_{1},\cdots,a_{n}|\forall a_{i}\in\mathcal{A}\}$ for $P$ in Eq. 1 to transit from $s_{0}$ to $\exists s_{n}\in S^{\star}$ in finite steps. Now, we explain each step in our planning pipeline to find a valid $\pi$ for $P$ and provide a more detailed explanation of the subgoal planner in the next section. An overview of our pipeline is also illustrated in Fig. 1.
@@ III-A Planning Formulation
For the robot to fully understand and interact with its environment, both semantics and geometry about the objects in the environment are required. We use a multi-modal LLM such as GPT-4o [^1] https://openai.com to simultaneously process both image and text prompts. By providing a color image captured by an RGBD camera along with the prompt, e.g., ”What objects are on the table? Tell me each of their appearance and spatial relationships.”, the LLM can describe the objects on the table, including their spatial relationships, positions, and appearance. Given the scene description, user-provided goal task, the domain PDDL, and an in-context example, the LLM generates a problem PDDL consisting of environment objects $\mathcal{O}$, the initial state $s_{0}$, and the goal state $S^{\star}$ to specify the planning problem $P$. We utilize one-shot prompting [21] by providing an example of problem PDDL generation to enhance the LLM’s responses.
We also employ a 2D open-vocabulary object detection model [22] to estimate the geometric information, specifically the bounding box of the target objects identified by the multi-modal LLM. These bounding boxes are essential for a robot manipulator to motion-plan their grasp poses.
@@ III-B Subgoal Generation
$P_{i}\equiv\langle\mathcal{S},\mathcal{A},\mathcal{T},s_{i},S^{\star}_{i+1}\rangle.$ (2)
We prompt the LLM with domain knowledge and a one-shot planning example along with the explanation of the steps for solving the problem and then ask the LLM to generate $\mathcal{G}$ by observing how the example problem is solved. For instance, in the Blocksworld-new domain, if the blocks are stacked in the order (on b1 b2)(on b2 b3), the reverse order stacking requires each of the three blocks to be unstacked with no objects on each block—(clear b1)(clear b2)(clear b3)—to rearrange them appropriately.
@@ IV-A Symbolic LLM Planner
When the size $|P_{i}|$ of $P_{i}$ is moderate, it is possible to use a symbolic planner to solve $P_{i}$ precisely. However, estimating $|P_{i}|$ is not easy. In theory, one can use a problem measure like MDL[4] to estimate it, but in practice, deriving the MDL for a challenging task is quite hard. Instead, one may estimate an MDL-like metric for $P_{i}$ by empirically measuring the planning time spent by running the MCTS LLM planner or a symbolic planner for a sampled $P_{i}$. If such an estimate is sufficiently high, we assume that $P_{i}$ is complex and resort to the MCTS LLM planner in the next section; otherwise, we use a symbolic planner.
@@ IV-B MCTS LLM Planner
When $|P_{i}|$ is high, using a symbolic planner to solve $P_{i}$ is impractical due to the high combinatorial search space. In this case, we use an MCTS planner combined with the LLM. As illustrated in Fig. 2, our MCTS LLM planner first samples $n_{s}$ plans for a sub-problem $P_{i}$ using an LLM (i.e., L-Policy), followed by building a state tree with the LLM-sampled plans, which serves as the reduced search space. The MCTS algorithm then searches this tree to identify an action sequence (i.e., a policy) that leads to a state satisfying the subgoal $S^{\star}_{i+1}$.
@@ IV-B1 Plan Sampling
Given the domain PDDL in Sec. III-A and a couple of in-context planning examples, the LLM generates $n_{s}$ best plans, $\{\pi^{1}_{i},\pi^{2}_{i},\cdots,\pi^{n_{s}}_{i}\}$, to achieve the subgoal in $P_{i}$. This is reminiscent of [5], but unlike [5], where they sample the entire problem $P$, we sample only for a sub-problem $P_{i}$, which is presumably more accurate. Also, the action weight is computed by summing the token log probabilities corresponding to the action generated by the LLM during plan sampling, reflecting the LLM’s confidence when generating the action [25]. Since a token’s log probability represents the token’s conditional probability given the preceding tokens, the action weight can be viewed as the conditional probability of the current action occurring, given the history of previous actions. This action weight will guide the rollout process in the MCTS.
@@ IV-B2 State Tree Generation
We generate a state tree $T_{i}$ for $P_{i}$ by coalescing the sampled $n_{s}$ plans where a node in $T_{i}$ is a state $s\in\mathcal{S}$ and an edge is an action $a\in\mathcal{A}$ connecting $s,s^{\prime}$ when $s^{\prime}=\mathcal{T}(s,a)$. $T_{i}$ bounds the search space for the MCTS to explore later, ensuring the search is constrained within the valid actions generated by the LLM. Moreover, we check if the preconditions of $a$, which is included in the domain PDDL, match the state $s$, then $a$ is valid and thus added to $T_{i}$; otherwise, subsequent actions are removed from $T_{i}$. This post-validity test is performed for each action in all the sampled plans.
@@ IV-B3 Monte Carlo Tree Search
1. Selection: Starting from the root node $s_{i}$, we recursively traverse $T_{i}$ by selecting the child node with the highest UCB1 score[26] from the set of visited nodes until we arrive at a node whose all child nodes are visited for the first time. Then, one of the child nodes is randomly selected, say $s_{r}$. If $s_{r}$ is included in the goal states $s_{r}\in S^{\star}_{i+1}$, the MCTS stops immediately and $s_{r}$ is traced back to $s_{i}$, thereby constructing a plan $\pi_{i}$ for $P_{i}$.
@@ V-A Experimental Setup
All task planning experiments were conducted on an Intel Core i9 CPU and NVIDIA RTX 6000 GPUs. We employed GPT-4o for the multi-modal LLM and used the fast downward planner as the symbolic PDDL planner. We conducted PDDL task planning experiments in three well-known IPC domains by modifying their problem complexities[27]: Barman-new, Blocksworld-new, and Gripper-new.
@@
In this domain, a robotic arm stacks $3\leq n\leq 10$ blocks, randomly divided into one to three stacks arranged on a table. The goal is to rearrange the blocks for each stack. Unlike the original Blocksworld domain, we increase the planning complexity by creating six block placement positions for interim workspace. As a result, the planner must also specify positions for placing the blocks rather than using a single on-table predicate as in the original domain.
@@
In the Gripper-new domain, four mobile robots move $2\leq n\leq 10$ balls to four different rooms from their initial location. We incorporate four multiple robots, making the planning process more complex in a multi-agent scenario, similar to [13]. The positions of the balls and robots in both the initial and goal states are random.

### v2

@@ Abstract
In robotic task planning, symbolic planners using rule-based representations like PDDL are effective but struggle with long-sequential tasks in complicated environments due to exponentially increasing search space. Meanwhile, LLM-based approaches, which are grounded in artificial neural networks, offer faster inference and commonsense reasoning but suffer from lower success rates. To address the limitations of the current symbolic (slow speed) or LLM-based approaches (low accuracy), we propose a novel neuro-symbolic task planner that decomposes complex tasks into subgoals using LLM and carries out task planning for each subgoal using either symbolic or MCTS-based LLM planners, depending on the subgoal complexity. This decomposition reduces planning time and improves success rates by narrowing the search space and enabling LLMs to focus on more manageable tasks. Our method significantly reduces planning time while maintaining high success rates across task planning domains, as well as real-world and simulated robotics environments. More details are available at http://graphics.ewha.ac.kr/LLMTAMP/.
@@ I INTRODUCTION
At a high level, the use of LLMs for task planning is divided into treating LLMs as a policy model (known as L-Policy) or as a world model (known as L-Model)[4]. L-Policy exploits the commonsense knowledge of LLMs to directly query proper policy for a given state, while L-Model utilizes LLMs as a simulation model of the world to query the state of the world as a result of executing an action or a policy. However, despite their strengths, LLMs suffer from token inefficiency and correction inefficiency [5], often generating hallucinated action sequences and failing on complex tasks [6]. To address the limitations of current LLM-based task planners, we propose a novel neuro-symbolic task planner that leverages LLMs as both L-Policy and L-Model to solve a long-sequential robotic task. Our planner is significantly faster than symbolic planners and more accurate than LLM-based planners.
An immediate issue in handling long-sequential tasks using LLMs is LLM’s token inefficiency since the planning descriptions involve a long and repetitive sequence of world and robotic states and a history of policies and their results. To circumvent this issue, we utilize LLMs as L-Model to generate a sequence of subgoals for a long-horizon task, effectively decomposing it into smaller and manageable sub-tasks. This goal decomposition also provides a useful side-effect to reduce the overall search space, yielding an accurate subgoal planner based on LLMs. Indeed, we use the Monte Carlo Tree Search (MCTS) algorithm while using LLMs as L-Policy to accurately solve each subgoal, reducing the correction inefficiency common in LLM-based planners.
Furthermore, if the original task is moderately complex, requiring a smaller minimum description length (MDL)[4] to solve the given problem, one can rely on a symbolic planner to solve the subgoals precisely while effectively avoiding the exponential growth of planning time.
@@
    Planning formulation: Given a planning goal in natural language description and domain knowledge, our task planner relies on PDDL to encode the problem descriptions. We also obtain the semantic and spatial relationships of target objects in the environment using a multimodal LLM, translated and encoded in problem PDDL.
@@
    Task planning: If the MDL is moderate, we rely on a symbolic planner to solve each subgoal; otherwise, we generate and expand a search tree and use the MCTS algorithm with L-Policy as a rollout policy to solve the subgoal. This subgoal planning is repeated for each sub-task, and the plans are combined to form the overall plan.
We conducted experiments across three task planning domains while varying the problem complexity. Compared to the state-of-the-art symbolic task planner like the Fast Downward planner [3], our approach significantly reduced planning time while maintaining an acceptable success rate. Additionally, we conducted experiments using dual robot manipulators and a robotic simulator to demonstrate the practical utility of our planner.
@@
The rest of this paper is organized as follows. In Sec. II, we review relevant work to task planning. In Sec. III, we outline the overall pipeline, and in Sec. IV, we explain the algorithms of both the symbolic subgoal planner and the LLM-based subgoal planner. In Sec. V, we present the task planning results and experiments in real and simulation robotics environments and conclude the paper and discuss future work in Sec. VI.
@@ II-B LLM-based Robot Task Planning
Recent studies have explored using LLMs for robot task planning by leveraging their real-world understanding. [9] combines language understanding with action grounding in real-world affordances, enabling robots to execute tasks based on their capabilities. Similarly, [10] introduced a prompting scheme that enables LLM to generate Python codes composed of robot action primitives, incorporating environmental state feedback. [11] fine-tuned multimodal LLMs to integrate physical grounding with visual inputs for task planning. TAMP has also been addressed using LLM by [12], enabling LLM as spatial relationship generators between environment objects. Additionally, some studies combined LLM-based high-level planning with reinforcement learning for low-level control [13]. However, these approaches’ common limitations are low success rates in solving long sequential tasks, limited multi-step reasoning, and weak failure recovery.
Recent Large Reasoning Models have shown high success rates across various PDDL domains without additional frameworks, their performance still declines as task complexity increases, with success rates approaching zero in highly constrained domains [14].
@@ II-C Hybird Task Planning
Recently, studies have been conducted on integrating LLMs with symbolic planning methods. [15] and [16] used LLMs to translate natural language problem descriptions into PDDL initial states and goals through few-shot prompting. However, these approaches struggle in real-world applications where problems are not presented in natural language. [17] combined LLMs with vision models to generate planning problem specifications based on real-world scenes, using re-prompting to correct specification errors. LLMs have also been used to solve PDDL problems. [18] showed that while LLMs can solve some non-trivial PDDL problems, they often fail on more complex tasks, though their outputs can guide heuristic planners. [19] improved this by generating Python functions for PDDL planning with automated debugging, while [20] introduced an iterative refinement framework using validator feedback. While these methods have improved success rates compared to LLM-only methods, they have been tested mostly on small-scale problems.
@@ II-D Integrating LLMs with Tree Search
Combining tree structures with LLM-generated actions has been explored in various studies. [21] samples possible next actions from the current state using an LLM and selects the best one via an LLM-evaluator, iterating with DFS or BFS. To improve token and runtime efficiency, [5] proposes sampling multiple plans at once to generate an action tree and selecting actions from the tree based on observations and histories. [22] integrates MCTS with LLMs for iterative state transitions in MDPs, and [4] employs LLMs as both L-Model and L-Policy within MCTS to solve large-scale POMDPs.
Our method also samples multiple plans at once using an LLM and applies MCTS, but unlike [5], ours relies on LLM-induced goal decomposition to generate multiple deterministic sub-problems. Unlike [4], our MCTS operates on a fixed tree for a sub-problem rather than the entire planning problem.
@@ III TASK PLANNING PIPELINE
$P\equiv\langle\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},s_{0},S^{\star}\rangle,$ (1)
where $\mathcal{S}$ is a finite set of fully observable states,$\mathcal{O}$ is environment objects, $\mathcal{A}$ is a finite set of possible actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is a deterministic state transition function, $s_{0}\in\mathcal{S}$ is an initial state, and $S^{\star}\subset\mathcal{S}$ is a set of goal states. Our planning objective is to find a policy $\pi=\{a_{1},\cdots,a_{n}|\forall a_{i}\in\mathcal{A}\}$ for $P$ in Eq. 1 to transit from $s_{0}$ to $\exists s_{n}\in S^{\star}$ in finite steps. Now, we explain each step in our planning pipeline to find a valid $\pi$ for $P$ and provide a more detailed explanation of the subgoal planner in the next section. An overview of our pipeline is also illustrated in Fig. 1.
@@ III-A Planning Formulation
For the robot to fully understand and interact with its environment, both semantics and geometry about the objects in the environment are required. We use a multimodal LLM such as GPT-4o [^1] https://openai.com to process image and text prompts simultaneously. By providing a color image captured by an RGBD camera along with the prompt, e.g., ”What objects are on the table? Tell me each of their appearance and spatial relationships.”, the LLM can describe the objects on the table, including their spatial relationships, positions, and appearance. Given the scene description, user-provided goal task, the domain PDDL, and an in-context example, the LLM generates a problem PDDL consisting of environment objects $\mathcal{O}$, the initial state $s_{0}$, and the goal state $S^{\star}$ to specify the planning problem $P$. We utilize one-shot prompting [23] by providing an example of problem PDDL generation to enhance the LLM’s responses.
We also employ a 2D open-vocabulary object detection model [24] to estimate the geometric information, specifically the bounding box of the target objects identified by the multimodal LLM. These bounding boxes are essential for a robot manipulator to motion-plan their grasp poses.
@@ III-B Subgoal Generation
$P_{i}\equiv\langle\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},s_{i},S^{\star}_{i+1}\rangle.$ (2)
We prompt the LLM with domain knowledge and a one-shot planning example along with the explanation of the steps for solving the problem and then ask the LLM to generate $\mathcal{G}$ by observing how the example problem is solved. For instance, in the Blocksworld-new domain, if the blocks are stacked in the order (on b1 b2)(on b2 b3)(on-table b3 t1), the reverse order stacking requires each of the three blocks to be unstacked with no objects on each block—(clear b1)(clear b2)(clear b3)(clear-table t1)—to rearrange them appropriately.
@@ IV-A Symbolic LLM Planner
When the size $|P_{i}|$ of $P_{i}$ is moderate, it is possible to use a symbolic planner rather than an LLM-based planner to solve $P_{i}$ precisely. However, estimating $|P_{i}|$ is not easy. In theory, one can use a problem measure like MDL [4] to estimate it, but in practice, deriving the MDL for a challenging task is quite hard. Instead, one may estimate an MDL-like metric for $P_{i}$ by empirically measuring the planning time spent by running the MCTS LLM planner or a symbolic planner for a sampled $P_{i}$. If such an estimate is sufficiently high, we assume that $P_{i}$ is complex and resort to the MCTS LLM planner in the next section; otherwise, we use a symbolic planner.
@@ IV-B MCTS LLM Planner
When $|P_{i}|$ is high, using a symbolic planner to solve $P_{i}$ is impractical due to the high combinatorial search space. In this case, we use an MCTS planner combined with the LLM. As illustrated in Fig. 2, our MCTS LLM planner first samples $n_{s}$ plans for a sub-problem $P_{i}$ using an LLM (i.e., L-Policy), then building a state tree with the LLM-sampled plans, which serves as the reduced search space. The MCTS algorithm then searches this tree to identify an action sequence (i.e., a policy) that leads to a state satisfying the subgoal $S^{\star}_{i+1}$.
@@ IV-B1 Plan Sampling
Given the domain PDDL in Sec. III-A and a few in-context planning examples, the LLM generates $n_{s}$ plans, $\{\pi^{1}_{i},\pi^{2}_{i},\cdots,\pi^{n_{s}}_{i}\}$ to achieve the subgoal in $P_{i}$. Unlike [5], which samples the entire problem $P$, we sample only for a sub-problem $P_{i}$, leading to presumably higher accuracy. Also, the action weight is computed as the sum of token log probabilities for each LLM-generated action, reflecting the LLM’s confidence when generating the action [27]. Since a token’s log probability represents its conditional probability given previous tokens, the action weight can be viewed as the conditional probability of the current action occurring, given the history of previous actions. This action weight will guide the rollout process in the MCTS.
@@ IV-B2 State Tree Generation
We generate a state tree $T_{i}$ for $P_{i}$ by coalescing the sampled $n_{s}$ plans where each node in $T_{i}$ represents a state $s\in\mathcal{S}$ and each edge corresponds to an action $a\in\mathcal{A}$ connecting $s,s^{\prime}$ when $s^{\prime}=\mathcal{T}(s,a)$. $T_{i}$ bounds the MCTS search space, ensuring the search is restricted to valid LLM-generated actions. Moreover, we verify whether the preconditions of $a$, as defined in the domain PDDL, hold for $s$. If valid, then $a$ is added to $T_{i}$; otherwise, subsequent actions are removed from $T_{i}$. This post-validity check is applied to every action in all sampled plans.
@@ IV-B3 Monte Carlo Tree Search
1. Selection: Starting from the root node $s_{i}$, we recursively traverse $T_{i}$ by selecting the child node with the highest UCB1 score[28] (with an exploration parameter of 1) from the set of visited nodes. This continues until reaching a node whose all child nodes are visited for the first time. Then, one of the child nodes is randomly selected, say $s_{r}$. If $s_{r}$ is included in the goal states $s_{r}\in S^{\star}_{i+1}$, the MCTS stops immediately and $s_{r}$ is traced back to $s_{i}$, thereby constructing a plan $\pi_{i}$ for $P_{i}$.
@@ V-A Experimental Setup
All task planning experiments were conducted on an Intel Core i9 CPU and NVIDIA RTX 6000 GPUs. We employed GPT-4o for the multimodal LLM with temperature 0.0 and used the Fast Downward planner as the symbolic PDDL planner. We conducted PDDL task planning experiments in three well-known IPC domains by modifying their problem complexities[29]: Barman-new, Blocksworld-new, and Gripper-new.
@@
In this domain, a robotic arm stacks $3\leq n\leq 10$ blocks, randomly divided into one to three stacks arranged on a table. The goal is to rearrange the blocks for each stack. Unlike the original Blocksworld domain, we increase the planning complexity by creating six block placement positions for the interim workspace. As a result, the planner must also specify positions for placing the blocks rather than using a single on-table predicate as in the original domain.
@@
In the Gripper-new domain, four robots move $2\leq n\leq 10$ balls to four different rooms from their initial location. We incorporate four multiple robots, making the planning process more complex in a multi-agent scenario, similar to [15]. The positions of the balls and robots in both the initial and goal states are random.